OpenStudy (pottersheep):

Pleaseeee help, Grade 12! If x^3 - 4x^2 +kx =0 what value of k would result in two equal roots?

OpenStudy (anonymous):

x(x^2-4x+k)=0 16-4k=0 so k=4

OpenStudy (pottersheep):

thank you so much

OpenStudy (anonymous):

np

OpenStudy (pottersheep):

sorry where did 16 -4k come from?

OpenStudy (anonymous):

b-sqrt(b^2-4ac)... that formula. so b^2-4ac=4^2-4k. as you can see, you can factor out an x to get a quadratic equation, use the quadratic formula to solve that equation and there you go :)

OpenStudy (pottersheep):

oooo thanks!